A polygon has 35 diagonals. How many sides does it have?
Answer by edjones(8007)

Let d=number of diagonals and n=number of sides
(n(n-3))/2=d
(n^2-3n)/2=35
.5n^2-1.5n-35=0
n^2-3n-70=0
(n-10)(n+7)=0
n=10
.
Ed